# World Models via Policy-Guided Trajectory Diffusion

### v2

@@ 5.3 Can the synthetic data produced by PolyGRAD be used to train performant policies?
To assess the performance of RL training in imagination for PolyGRAD (Algorithm 3), we compare the rewards obtained against model-free on-policy RL algorithms (PPO (Schulman et al., 2017), TRPO (Schulman et al., 2015), and A2C (Mnih et al., 2016)). We compare against on-policy model-free RL algorithms because we perform imagined on-policy training in the PolyGRAD world model. However, note that off-policy model-free RL algorithms obtain stronger performance for these environments (Haarnoja et al., 2018). To compare against a state-of-the-art world model-based agent, we also compare against Dreamer-v3 (Hafner et al., 2023). The plots in Figure 6 show that PolyGRAD outperforms on-policy model-free algorithms. This is likely because PolyGRAD leverages all of the training data available to train the world model and generate large quantities of synthetic training data. On the other hand, on-policy model-free algorithms do not reuse training data.
@@ 5.4 Which implementation details influence the performance of PolyGRAD?
A detailed description of each of these modifications is provided in Appendix C.3. Figure 7 shows the reward curves for each variant of the algorithm. We observe that Random Actions obtains the poorest performance. This is unsurprising as we use on-policy RL to optimise the policy, and therefore randomly sampled off-policy actions result in incorrect policy gradient updates. Policy Sampling also performs very poorly. This method also fails to generate on-policy trajectories due to the following reasoning. Imagine that we sample an entire sequence of actions, conditioned on the current state sequence. Then, conditioned on these new actions, a new prediction of the state sequence is generated at the next diffusion step. However, the policy has a different action distribution at the new state sequence. Thus, the final trajectory of states and actions produced is not on-policy. We also observed that Policy Sampling results in worse prediction errors. This is likely because a diffusion model, which gradually refines its predictions, is ill-suited to making an accurate prediction when the conditioning is completely changed at each diffusion step. Meaningful policies are still learnt when we ablate the action clipping (No Clipping), however the performance is worse, especially for Walker2d. The plots in Appendix B.3 show that without the action clipping, the action distributions produced are more heavy-tailed. Conditioning the policy on the noisy states during diffusion, rather than the denoised prediction (Noisy State Conditioning), also results in worse performance. Additionally updating the states during diffusion according to Equation 8 results in comparable performance to the base PolyGRAD algorithm. However, this modification increases the computation time required to about 80 hours for 1M environment steps (compared to 54 hours for the base algorithm), as it requires backpropagation through the policy at each diffusion step.
@@
Figure 7: Reward curves for PolyGRAD, and ablations and modifications to PolyGRAD. The shaded region indicates standard deviation across 4 seeds for each variant and 5 seeds for the base PolyGRAD algorithm.
@@ 6 Discussion
To improve the stability issues mentioned above, we plan to investigate alternative algorithms for diffusing on-policy trajectories, with the aim of finding an algorithm that more reliably generates the correct action distribution. We would like to extend PolyGRAD to image-based environments by utilising latent diffusion (Rombach et al., 2022). We would also like to test PolyGRAD on non-Markovian environments. PolyGRAD is well-suited to non-Markovian predictions due to the fact that it is trained to predict entire trajectories, rather than making single step predictions under a Markovian assumption. Another direction for future work is to perform a theoretical analysis of the convergence properties of PolyGRAD to better understand if there are conditions under which it is guaranteed to converge to on-policy trajectories. Finally, we would like to perform more experiments to investigate whether there are situations in which PolyGRAD obtains better prediction errors than the strongest baseline, Autoregressive Diffusion. We hypothesise that PolyGRAD may be more robust when trained on small datasets, where the predictions of single-step autoregressive models may be prone to quickly leaving the data distribution, resulting in erroneous predictions.
@@ Conclusion
We have presented PolyGRAD, the first scalable world model that can generate on-policy trajectories without autoregressive sampling. To enable this, instead of sampling actions directly from the policy, PolyGRAD gradually updates the trajectory of states and actions using diffusion. We have shown that PolyGRAD generates a good approximation to the correct on-policy action distribution, and therefore enables on-policy RL in imagination. PolyGRAD achieves strong performance in terms of trajectory prediction accuracy compared to state-of-the-art autoregressive baselines. We believe that PolyGRAD introduces a promising paradigm for world modelling, with many possible directions for extensions and improvements in future work.
@@ B.1 Prediction Error Plots with Transformer and MLP Denoising Networks
Figure 8: Plots of mean squared error (MSE) of predicted states vs ground truth states for PolyGRAD using either a transformer or residual MLP denoising network. Shaded regions indicate standard deviations over 4 seeds.
B.2 Extended Plots of PolyGRAD Action Distributions for All Environments
Figures 9--11 show plots of action distributions produced by PolyGRAD in each of the MuJoCo environments with $h=10$. Blue line illustrates the distribution of $a-\mu_{\phi}(s)$ for a batch of synthetic data. Each subplot is for a policy with a different entropy level. Dashed black line indicates the action distribution output by the policy.
@@ B.2 Extended Plots of PolyGRAD Action Distributions for All Environments
B.3 PolyGRAD Action Distributions with Action Clipping Ablated
@@ C.1 Experimental Setup
C.2 Baselines
@@ C.4 Computational Requirements
UhEUgAAAAsAAAAOCAYAAAD5YeaVAAAAAXNSR0IArs4c6QAAAAZiS0dEAP8A/wD/oL2nkwAAAAlwSFlzAAALEwAACxMBAJqcGAAAAAd0SU1FB9wKExQZLWTEaOUAAAAddEVYdENvbW1lbnQAQ3JlYXRlZCB3aXRoIFRoZSBHSU1Q72QlbgAAAdpJREFUKM9tkL+L2nAARz9fPZNCKFapUn8kyI0e4iRHSR1Kb8ng0lJw6FYHFwv2LwhOpcWxTjeUunYqOmqd6hEoRDhtDWdA8ApRYsSUCDHNt5ul13vz4w0vWCgUnnEc975arX6ORqN3VqtVZbfbTQC4uEHANM3jSqXymFI6yWazP2KxWAXAL9zCUa1Wy2tXVxheKA9YNoR8Pt+aTqe4FVVVvz05O6MBhqUIBGk8Hn8HAOVy+T+XLJfLS4ZhTiRJgqIoVBRFIoric47jPnmeB1mW/9rr9ZpSSn3Lsmir1fJZlqWlUonKsvwWwD8ymc/nXwVBeLjf7xEKhdBut9Hr9WgmkyGEkJwsy5eHG5vN5g0AKIoCAEgkEkin0wQAfN9/cXPdheu6P33fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v3

@@ 5.3 Can the synthetic data produced by PolyGRAD be used to train performant policies?
To assess the performance of RL training in imagination for PolyGRAD (Algorithm 3), we compare the rewards obtained against model-free on-policy RL algorithms (PPO (Schulman et al., 2017), TRPO (Schulman et al., 2015), and A2C (Mnih et al., 2016)). We compare against on-policy model-free RL algorithms because we perform imagined on-policy training in the PolyGRAD world model. However, note that off-policy model-free RL algorithms obtain stronger performance for these environments (Haarnoja et al., 2018). To compare against a state-of-the-art world model-based agent, we also compare against Dreamer-v3 (Hafner et al., 2023). Reward curves for each algorithm are shown in Figure 5.3. Figure 7 provides 95% confidence intervals of the interquartile mean of the normalised final performance aggregated across all three environments (see Appendix C.2 for details). Figure 7 shows that PolyGRAD outperforms on-policy model-free algorithms. This is likely because PolyGRAD leverages all of the training data available to train the world model and generate large quantities of synthetic training data. On the other hand, on-policy model-free algorithms do not reuse training data. Confidence intervals in Appendix B.2 show that PolyGRAD significantly outperforms model-free algorithms in terms of mean performance in addition to interquartile mean.
@@
Figure 7: 95% CIs of interquartile mean of normalised final performance aggregated across envs.
@@ 5.4 Which implementation details influence the performance of PolyGRAD?
A detailed description of each of these modifications is provided in Appendix C.4. Figure 5.4 shows the reward curves for each variant of the algorithm, and Figure 9 shows 95% confidence intervals for the interquartile mean of the final normalised performance across all environments. We observe that Random Actions obtains the poorest performance. This is unsurprising as we use on-policy RL to optimise the policy, and therefore randomly sampled off-policy actions result in incorrect policy gradient updates. Policy Sampling also performs very poorly. This method also fails to generate on-policy trajectories due to the following reasoning. Imagine that we sample an entire sequence of actions, conditioned on the current state sequence. Then, conditioned on these new actions, a new prediction of the state sequence is generated at the next diffusion step. However, the policy has a different action distribution at the new state sequence. Thus, the final trajectory of states and actions produced is not on-policy. We also observed that Policy Sampling results in worse prediction errors. This is likely because a diffusion model, which gradually refines its predictions, is ill-suited to making an accurate prediction when the conditioning is completely changed at each diffusion step. Meaningful policies are still learnt when we ablate the action clipping (No Clipping), however the performance is worse, especially for Walker2d. The plots in Appendix B.4 show that without the action clipping, the action distributions produced are more heavy-tailed. Conditioning the policy on the noisy states during diffusion, rather than the denoised prediction (Noisy State Conditioning), also results in worse performance. Additionally updating the states during diffusion according to Equation 8 results in comparable performance to the base PolyGRAD algorithm. However, this modification increases the computation time required to about 80 hours for 1M environment steps (compared to 54 hours for the base algorithm), as it requires backpropagation through the policy at each diffusion step.
@@
Figure 8: Reward curves for PolyGRAD, and ablations and modifications to PolyGRAD. Shaded regions indicate standard deviation across 5 seeds.
Figure 9: 95% CIs of interquartile mean of normalised final performance aggregated across envs.
@@ 6 Discussion
To improve the stability issues mentioned above, we plan to investigate alternative algorithms for diffusing on-policy trajectories, with the aim of finding an algorithm that more reliably generates the correct action distribution. We would like to scale PolyGRAD to more complex environments such as image-based environments by utilising latent diffusion (Rombach et al., 2022). We would also like to test PolyGRAD on non-Markovian environments. PolyGRAD is well-suited to non-Markovian predictions due to the fact that it is trained to predict entire trajectories, rather than making single step predictions under a Markovian assumption. Another direction for future work is to perform a theoretical analysis of the convergence properties of PolyGRAD to better understand if there are conditions under which it is guaranteed to converge to on-policy trajectories.
Finally, and perhaps most importantly, we would like to investigate whether there are situations in which PolyGRAD obtains better prediction errors at long horizons compared to existing autoregressive baselines. The results in Section 5.2 show that while PolyGRAD obtains strong performance when trained and evaluated on short trajectories ($h=10$), performance deteriorates relative to baselines for longer trajectories. We hypothesise that PolyGRAD may be more robust than autoregressive models when trained on small datasets, where the predictions of single-step autoregressive models may be prone to quickly leaving the data distribution, resulting in erroneous predictions. Another line of investigation could be to train the denoising network on short trajectories, and use it to generate longer trajectories. Perhaps this may result in stronger performance and generalisation to longer trajectories. A final approach could be to investigate whether latent diffusion is better suited to accurate generation of long trajectories.
@@ Conclusion
We have presented PolyGRAD, the first world model that can generate on-policy trajectories without autoregressive sampling. To enable this, instead of sampling actions directly from the policy, PolyGRAD gradually updates the trajectory of states and actions using diffusion. We have shown that PolyGRAD generates a good approximation to the correct on-policy action distribution, and therefore enables on-policy RL in imagination. PolyGRAD achieves solid performance in terms of trajectory prediction accuracy compared to state-of-the-art autoregressive baselines. Thus, PolyGRAD introduces a promising new paradigm for world modelling, with many possible directions for extensions and improvements in future work.
@@ B.1 Prediction Error Plots with Transformer and MLP Denoising Networks
Figure 10: Plots of mean squared error (MSE) of predicted states vs ground truth states for PolyGRAD using either a transformer or residual MLP denoising network. Shaded regions indicate standard deviations over 5 seeds.
B.2 Confidence Intervals for Mean Performance
In the main body of the paper, we reported 95% confidence intervals for the normalised interquartile mean (IQM) generated using the rliable framework (Figures 7 and 9). To provide a more complete picture, here we provide equivalent confidence intervals for the mean, rather than the IQM. We observe that the confidence intervals for mean are similar to those for IQM. This indicates that there are not many outliers in the results, as the mean is effected by outliers while IQM is not.
The intervals in Figure 11 compare PolyGRAD against other model-based and model-free algorithms. We observe that PolyGRAD outperforms the model-free algorithms for mean performance, but does not perform as well as Dreamer-v3. Figure 12 compares the base PolyGRAD algorithm against each ablation and variant of the algorithm. PolyGRAD significatly outperforms each variant/ablation except for Add State Update.
Figure 11: 95% confidence intervals of mean of normalised final performance aggregated across environments.
Figure 12: 95% confidence intervals of mean of normalised final performance aggregated across environments for PolyGRAD and each variant.
B.3 Extended Plots of PolyGRAD Action Distributions for All Environments
Figures 13--15 show plots of action distributions produced by PolyGRAD in each of the MuJoCo environments with $h=10$. Blue line illustrates the distribution of $a-\mu_{\phi}(s)$ for a batch of synthetic data. Each subplot is for a policy with a different entropy level. Dashed black line indicates the action distribution output by the policy.
@@ B.3 Extended Plots of PolyGRAD Action Distributions for All Environments
B.4 PolyGRAD Action Distributions with Action Clipping Ablated
@@ C.1 Experimental Setup
C.2 Results Processing
Smoothing of Reward Curves
To generate the reward curves in Figures 5.3 and 5.4, the average return over 10 episodes of the current policy is evaluated after every 10,000 environment steps. To generate the plots we perform smoothing by computing a moving average with a window width of 10 evaluations.
Confidence Intervals
Figures 7 and 9 present 95% confidence intervals of the interquartile mean (IQM) of the normalised final performance aggregated across all three environments. Figures 11 and 12 are equivalent plots for the mean rather than IQM. To generate the confidence intervals, we use the rliable framework (Agarwal et al., 2021) which is designed to compute confidence intervals with limited seeds by aggregating the runs across all environments, and generating the confidence intervals using stratified bootstrapping. IQM is the performance on the middle 50% of combined runs, and is the metric recommended by rliable as it is robust to outliers (unlike the mean) and is more stastically efficient than the median.
To generate the confidence intervals, we first aggregate the final smoothed rewards obtained at the maximum number of environment steps for each run of each algorithm in Figures 5.3 and 5.4. Following Agarwal et al. (2021), we first normalise the rewards in each environment. We normalise the rewards by dividing by the best average final total reward obtained by any algorithm for each environment. We then use the normalised final rewards for each run to compute the aggregated IQM confidence intervals using the rliable package with 50,000 bootstrapping repeats.
C.3 Baselines